OpenStudy (anonymous):

find the horizontal asymptote of the graph of y=-2x^6+5x+8/8x^6+6x+5

OpenStudy (phi):

ignore all terms except the highest order term in the top and bottom then find the ratio of those two terms.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

@thesecretstoday?

OpenStudy (anonymous):

Divide up and down by x^6 to obtain \[\Large \frac{-2 x^6+5 x+8}{8 x^6+6 x+5}=\frac{-2 x^6+5 x+8}{\frac{x^6 \left(8 x^6+6 x+5\right)}{x^6}}=\frac{\frac{8}{x^6}+\frac{5}{x^5}-2}{\frac{5}{x^6}+\frac{6}{x^5} +8} \] If x goes to infinity, what is the limit of the ratio?

OpenStudy (anonymous):

The limit goes to \[ \frac {0 + 0 -2}{0+0+8}=-\frac 2 8 =-\frac 1 4 \] so y= 1/4 is your vertical asymptote.

OpenStudy (anonymous):

uhhhh sooo is the horizontal the same thing?

OpenStudy (anonymous):

Yes